OpenStudy (anonymous):

Find the unit rate in feet per second 180miles/3hours Show your work

OpenStudy (saifoo.khan):

60

OpenStudy (anonymous):

Nope

OpenStudy (anonymous):

180/3=60 miles/hour =60*6076feet/60*60sec =6076/60 =60. feet/sec